OpenStudy (anonymous):

FAN & MEDAL <3 f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (anonymous):

just go through the answer choices and plug in each number for n

OpenStudy (superdavesuper):

u got f(1) n f(2) already...by using f(n) = f(n - 1) + f(n - 2), u can calculate f(3) n f(4)

OpenStudy (anonymous):

i dont kno what to do , or how to plug this in

OpenStudy (anonymous):

i dont see the f(3) or f(4) btw

OpenStudy (superdavesuper):

I will do f(3) here....f(3) = f(3-1) + f(3-2)

OpenStudy (superdavesuper):

f(3-1)=f(2) f(3-2)=f(1) so....

OpenStudy (anonymous):

that's looking like a foreign language right now idk what is going on

OpenStudy (superdavesuper):

it is a mix of english n numbers ;) OK tell me this what is f(1)=? and f(2)=?

OpenStudy (anonymous):

f(1)=1 ? and f(2)=2 ?

OpenStudy (superdavesuper):

OK then f(3) = f(3-1) + f(3-2) = f(2) + f(1)

OpenStudy (anonymous):

im not sure where that f("3") is coming from but ok f(2) + f(1) then i ?

OpenStudy (superdavesuper):

think u need to read ur question again....it is asking u to list the series of f(n) right?

OpenStudy (superdavesuper):

f(n) where n=1,2,3,4,5....

OpenStudy (anonymous):

I never understood it to begin with , but yes your right "for n" it says

OpenStudy (anonymous):

so the answer is 12345 ?

OpenStudy (superdavesuper):

OK read it again plz - u cannot answer a Q w/o understanding wat its asking in the 1st place

OpenStudy (superdavesuper):

no it is not 12345: that would be too simple! lol

OpenStudy (anonymous):

I re-read the question and no I cant understand it

OpenStudy (superdavesuper):

i can give u the ans but it wont do u much good...u need to ask ur teacher to explain it better ;)

OpenStudy (anonymous):

I would appreciate the answer, and I will ask my teacher to teach this to me

OpenStudy (superdavesuper):

Ans is 1, 2, 3, 5, 8 good luck :)

OpenStudy (anonymous):

thank you .

OpenStudy (superdavesuper):

welcome